You are here: Foswiki>Commissioning Web>Com20080618 (19 Jun 2008, AndrewRakich)Edit Attach
--++ Commissioning Notes from 17th-18th June 2008 MST

Observers: D. Miller, A. Rakich (in Tucson).
OSA: A. Ceranski

Summary

First part of the night sorting out a bug on offset pointing, then confirming that it seemed to work well. Second part of night taking field aberration data after puting a large rotation into M2. Early results are positive but need to analyze data.

Night time Activities

(Times are in MST)

8:30 guiding on-axis on BS9158 , having problem related to issue 1447. Images in /temp/guiderimage029950-030025, The WFSC image that has the PA, RA Dec etc. in the header is 080618w/wfsc000013_dd. fits

8:35 On the TO station, 101 and 102 TCS died; on 107 GCS died; all the TO station GUIs died. Steve has no idea why this happened but this was related to a change of directory that Steve had just made related to the offset pointing stuff.

8:55 /temp/guiderimage0030200-030250, WFSC image capturing RA DEC etc. 080618w/wfsc000013_dd. fits.

10:20 low level subsytem controlling el drives was getting bad data, fixed by reset.

10:22 Going to star BS9158 #4 for Michelle's test for offset pointing.

10:35 OAC returned error 201 for reasonable coordinates: (-0.6, 517). Don't know what error 201 is. RPC communication error reported.

10:46 We are having trouble talking to OAC server more often while doing offsetting than in normal operation. 2 times failed for 4 offsets.

12:16 repeat of failure mentioned at 10:35 above when we try offset of (20, -20) BS9158 #6. Offsets of (5,-5) work, (20,-20) failed twice then worked. 080618w/wfsc000013_dd captures RA, Dec etc. in the header.

12:22 Michele declares partial victory and leaves.

12:38 go to BS1977 on axis and collimate with ActO. Guiding behaving poorly, have to keep manually updating with PTmodify. wfsc 20-23

12:46 goto BS9177#32 taking data wfsc 26-29

12:58 back on-axis to collimate wfsc and can't find star. have no explanation.

1:26 Offsets for WET star IE -76 CA -23, now driving in TIp +45", and find that as expected this needs a CA of -45"

1:44 back on BS9177 starting again and recollimating with ActO NOT bacuae Tom had reasonable request to go to horizon for a few minutes and as we were just about to start, we might as well let him.

2:00 back on BS9177 on axis to collimate. wfsc35-44. Going to star # 57, taking data 46-51

2:30 Guiding is just not working out at all, this is the worst it has been since we started. Doug is guiding by hand so we can take WFS data. Later will produce some failing guiding examples for Torsten and Michele.

2:32 going on axis and recollimating with active optics wfsc53-55

2:40 going to #19 , hit rotator dead band. going to #51, WFSC 63-65; going back on axis for acto recollimate wfsc 67-72

2:55 guiding seems to be better with longer exposure. going to # 25. taking data wfsc 74-79

3: 20 going back on-axis re-collimate with ActO wfsc 84-90 problems with guiding again improved by setting exposure to 900 ms, giving signal ~ 1500, background ~ 1300.

3:42 going to satar 52 wfsc 92-94 ; going back on axis recollimate wfsc 96, 98-99

3:54 going to #34 WFSC 101-103

3:56 going back on axis to recollimate wfsc 106-107

4:04 going to #35 wfsc 11-12, 114-116 but highly suspect data now (first blush of dawn spoiling our images

4:20 ever the optomist, back on-axis collimating: wfsc 118-119

4:22 going to #14 getting the error "did not receive full ccd readout much more frequently from GCS tonight than previously.

.

-- AndrewRakich - 18 Jun 2008
Topic revision: r1 - 19 Jun 2008, AndrewRakich
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback